OpenStudy (anonymous):

Ultraviolet rays, which cause sub burn by penetrating only the top level of skin, have a wavelength about 1000 times the wavelength of an X-ray find the wavelength of the ultraviolet rays. X-ray wavelength = 10^-10 please show me how to find the wavelength of the ulraviolet ray

OpenStudy (anonymous):

also this too! The wavelength of visible lights are between 3.8 X 10^-7 m and 7.6 x 10^-7 are these wavelengths longer or shorter than those of vltraviolet rays?

OpenStudy (anonymous):

UV wavelength=1000* X-ray wavelength X-ray wavelength=10^-10 so, UV wavelength= 1000*10^-10 =(10^3)*(10^-10)=10^-7

OpenStudy (anonymous):

visible light wavelength longer than UV

OpenStudy (anonymous):

why is it longer? explain

OpenStudy (anonymous):

wave length \[\lambda\] has inverse relation to the frequency f, you can check the attached file have a better sight

OpenStudy (anonymous):